OpenStudy (anonymous):

A mountain climber jumps a 1.9-m-wide crevasse by leaping horizontally with a speed of 9.0 m/s. If the climber's direction of motion on landing is −45°, what is the height difference between the two sides of the crevasse?

OpenStudy (anonymous):

ok well the answer i've got doesn't use the angle provided but im pretty sure it should work all the same. so the variables are: Horizontal: u=9 Vertical: u=0 s=1.9 s=? a=0 a=-9.8 t=? t= the same as the horizontal therefore, s=ut 1.9=(9).t t=0.21111 seconds this means that s(vertical)= 1/2at^2 = (0.5).(-9.8).(o.2111)^2 =-0.4368m =0.4368 meters

OpenStudy (stormfire1):

First, let's see what we know: \[Vix=9.0 m/s\]\[Vi y=0 m/s\]\[Ax=0 m/s^2 \]\[Ay=-9.8 m/s^2 \]The length of the crevasse doesn't matter in this particular question. Since the jumper lands at a -45 degree angle the final Vy must have the same magnitude as the final Vx. Now you have enough information to solve the problem using : \[Vf^2 = V0^2+2ad\]Plugging in values: \[(-9.0m/s)^2 =0m/s + 2(-9.8 m/s^2)d\]Solving for d you get a distance (height difference) of -4.13m.